OpenStudy (anonymous):

Write a quadratic equation with the solution of -6

OpenStudy (saifoo.khan):

x = -6 (x+6)(x+6) = 0 x^2 + 6x + 6x + 36 = 0 x^2 + 12x + 36 = 0

OpenStudy (karatechopper):

did i tell you my brother was smart? dang he is like the bomb!

OpenStudy (saifoo.khan):

lol

OpenStudy (karatechopper):

its true right?!

OpenStudy (anonymous):

if Saifoo is your brother then yes he is very smart...

OpenStudy (karatechopper):

yes he is brother!